# Multi-layer diffusion model of photovoltaic installations

### v1

@@ 4 Conclusions
In this research, we proposed a new agent-based model that describes both opinion formation and diffusion of photovoltaic panels. Although, it is based on the well-known model of binary opinion dynamics, the $q$-voter model, we extended it by adding a second agent attribute (adoption state) and rewrote it to a 2-layer network structure. To generalize the model from a single-layer, we considered two variants. First, when $q$-voter’s social influence is required on both layers of the network to impact one’s opinion (AND variant). Second, where influence from just one layer is sufficient (OR variant). We investigated the model, the effect of parameters on exit states and exit times, using both Monte Carlo computer simulations and Mean-Field Approximation. For the approximation’s results, we used analytical methods wherever it was possible, and numerical where it was not.
@@
Let us begin with our main goal, which was to build a diffusion of innovation model that can replicate the phenomenon from the real world. Its process of adoption (for a certain range of parameter values) clearly resembles the growth of installed power in the data (Fig.1). However, it does not reduce the dynamics to a one-dimensional mechanism, as classical models did ([13]), but combines the adoption process with opinion formation. Consequently, diffusion of innovation and opinion dynamics are dependent on each other, with parameters only impacting one of these directly, influencing the other as well.
@@
More specifically, within the model there exist three possible exit states: unadopted system, adopted and finally, disordered one. The system transitions from the first, through the second, to the third, as the probability of independence $p$ increases. Broadly speaking, independence is initially essential for the adoption process to take off. However, as the diffusion accelerates, independence hinders innovation and ultimately prevents full adoption. Except for $a_{1}=0$, the probability of adoption has no impact on the exit states, only on exit times. When $a_{1}=0$, the system can never leave the unadopted state, should we set it as the initial one. What impacts the exit states, however, is the relation between $a_{1}$ and probability of loosing adoption $a_{2}$. As $a_{2}$ becomes less and less of $a_{1}$, critical values of $p$ required to reach the adopted state decrease, while those required for the disordered one increase. Basically, the greater the difference between people’s willingness to install solar panels versus their willingness to get rid of them, the better for the innovation. This difference not only amplifies independence in the initial phase of adoption, but dilutes it in the final phase as well.
@@
Regarding the studied variants, there is a significant difference between the two, AND and OR. In AND variant, other exit states are achieved more easily than in OR variant, i.e. lower values of $p$ are sufficient, but the transitions between these exit states are more pronounced. This is because, in AND variant the initial negative consensus is more easily disrupted than in OR variant, but so is the later positive consensus.
@@
Finally, it should be noted that MFA approximates the simulation results fairly well. Although, in the simulation version the network consists of Square Lattice and 2-dimensional Watts-Strogatz graph, structures pretty far from complete graph. Still, there are visible differences. Critical values of $p$ to reach various exit states differ slightly, but the general picture remains.

### v2

@@ 4 Conclusions
In this research, we proposed a new agent-based model that describes both opinion formation and diffusion of photovoltaic panels. Although, it is based on the well-known model of binary opinion dynamics, the $q$-voter model, we extended it by adding a second agent attribute (adoption state) and rewrote it to a two-layer network structure. To generalize the model from a single-layer, we considered two variants. First, when the $q$-voter’s social influence is required on both layers of the network to impact one’s opinion (the AND variant). Second, where influence from just one layer is sufficient (the OR variant). We investigated the model, the effect of parameters on stationary states and times to reach them, using both Monte Carlo computer simulations and Mean-Field Approximation. For the approximation’s results, we used analytical methods wherever it was possible, and numerical ones otherwise.
@@
Let us begin with our main goal, which was to build a diffusion of innovation model that does not reduce the dynamics to a one-dimensional mechanism, as classical models do [13], but combines the adoption process with opinion formation. Consequently, diffusion of innovation and opinion dynamics are dependent on each other, with parameters only impacting one of these directly, influencing the other as well.
@@
More specifically, within the model there exist three possible final states: unadopted system, adopted and finally, disordered one. The system transitions from the first, through the second, to the third, as the probability of independence $p$ increases. Broadly speaking, independence is initially essential for the adoption process to take off. However, as the diffusion accelerates, independence hinders innovation and ultimately prevents full adoption. Except for $a_{1}=0$, the probability of adoption has no impact on the stationary states, only on times to reach them. What impacts stationary states, however, is the relation between $a_{1}$ and probability of loosing adoption $a_{2}$. As $a_{2}$ becomes less and less of $a_{1}$, critical values of $p$ required to reach the adopted state decrease, while those required for the disordered one increase. Basically, the greater the difference between people’s willingness to install solar panels versus their willingness to get rid of them, the better for the innovation. This difference not only amplifies independence in the initial phase of adoption, but dilutes it in the final phase as well.
@@
Regarding the studied variants, there is a significant difference between the two, AND and OR. In the AND variant, other states are achieved more easily than in the OR variant, i.e. lower values of $p$ are sufficient, but the transitions between these states are more pronounced. This is because, in the AND variant the initial negative consensus is more easily disrupted than in the OR variant, but so is the later positive consensus.
@@
Finally, it should be noted that MFA approximates the simulation results fairly well. Although, in the simulation version the network consists of Square Lattice and two-dimensional Watts-Strogatz graph, structures pretty far from complete graph. Still, there are visible differences. Critical values of $p$ to reach various states differ slightly, but the general picture remains.
Appendix
In the basic research, only $c_{A}(0)=0$, $c_{S}(0)=0$ initial state was considered. However, one could investigate others as well. Here, we examine, what happens if initially a small fraction of agents possess positive opinions and adoption states, $A(0)=+1$, $S(0)=+1$. These agents could be chosen randomly or determined by some centrality measure. We compare time trajectories for the basic case (no such agents), with a number of such agents chosen randomly and the same number determined by highest degree on the second layer (degree centrality).
Figure 8: 10 simulated time trajectories of $c_{A}$ (red) and $c_{S}$ (blue) for $c_{A}(0)=0$, $c_{S}(0)=0$ (left), 100 agents with $A(0)=+1$, $S(0)=+1$ chosen randomly (middle) and 100 agents with $A(0)=+1$, $S(0)=+1$ determined by highest degree on the second layer (number of neighbors, right). The AND variant (top) and the OR variant (bottom). First layer – SL(N,1), second layer – WS2D(N,1,0.2), size $N=2500$ and $a_{2}=0.5a_{1}$ in all cases.
Results are shown in Fig. 8. In AND variant there are no visible changes, as such agents easily loose either their positive opinion or adoption state when their numbers are low. Then, with the AND rule in play, they have negligible impact on the whole system. In the OR variant however, there are significant differences. Existence of such agents visibly decreases the time needed for the system to reach the adopted state. Method of their choice bears a lesser impact. This is due to the fact, that the second layer is described by the two-dimensional Watts-Strogatz graph, WS2D($N=2500$, $m=1$, $\beta=0.2$). Distribution of degrees of such a random graph corresponds to the binomial (for small networks, $N=10^{2}$) or the Poisson distribution (for larger networks, $N\geq 10^{3}$) [51]. Therefore, the network lacks agents with degrees greatly exceeding the average, who could bear an impact.
Acknowledgements
T. W. was supported by the Ministry of Science and Higher Education, Poland, within the “Diamond Grant” Program through grant no. DI2019 0008 49.